OpenStudy (anonymous):

Solve the equation cos(θ + 60◦) = 2 sin θ giving all solutions in the interval 0◦ ≤ θ ≤ 360◦.

OpenStudy (anonymous):

use addition angle formula for first one is probably a good start

OpenStudy (anonymous):

Not sure what that is... can someone show it step by step? I'm not familiar with these equations?

OpenStudy (anonymous):

\[\cos(x+60)=\cos(x)\cos(60)-\sin(x)\sin(60)\] \[=\frac{1}{2}\cos(x)-\frac{\sqrt{3}}{2}\sin(x)\] hmmm maybe this was not such a good start

OpenStudy (anonymous):

wow, ok i am stuck. this is how wolfram does it, and i'll be darned if i see an easy method http://www.wolframalpha.com/input/?i=cos%28x%2Bpi%2F3%29%3D2sin%28x%29 what class is this for?

OpenStudy (anonymous):

It's for A-levels, pure mathematics 3. I'll check it out and tell you if I need help.

OpenStudy (anonymous):

I have a stepped solution, but am not sure what it all means.. If I give you that, maybe you can help me explain each step?

OpenStudy (anonymous):

ok i have an idea, but the solutution would require a calculator

OpenStudy (anonymous):

first we can write \[=\frac{1}{2}\cos(x)-\frac{\sqrt{3}}{2}\sin(x)=2\sin(x)\] then \[\cos(x)-\sqrt{3}\sin(x)=4\sin(x)\] \[\cos(x)=(4-\sqrt{3})\sin(x)\] \[\frac{\cos(x)}{\sin(x)}=4-\sqrt{3}\] or \[\frac{\sin(x)}{\cos(x)}=\frac{1}{4-\sqrt{3}}\]

OpenStudy (anonymous):

then we can take the inverse tangent to get a numerical answer

OpenStudy (anonymous):

if you have a better way please let me know, i am curious

OpenStudy (anonymous):

I'm lost.. Sorry! I'm not at all good with these things... if you can explain from the easiest possible step, I'd be much obliged! :)

OpenStudy (anonymous):

well i don't really know, that is the problem

OpenStudy (anonymous):

3 Attempt use of cos(A + B) formula to obtain an equation in cos θ and sin θ M1 Use trig formula to obtain an equation in tan θ (or cos θ, sin θ or cot θ) M1 Obtain tan θ = 1/(4 + 3 ) or equivalent (or find cos θ, sin θ or cot θ) A1 Obtain answer θ = 9.9° A1 Obtain θ = 189.9°, and no others in the given interval [Ignore answers outside the given interval. Treat answers in radians as a misread (0.173, 3.31).] This is what it says... not sure what it all means?

OpenStudy (anonymous):

first i used the addtion angle formula

OpenStudy (anonymous):

ok just what it says

OpenStudy (anonymous):

Can you explain this to me?

OpenStudy (anonymous):

yes. first we use this \[\cos(a+b)=\cos(a)\cos(b)-\sin(a)\sin(b)\] on the left hand side we have \[\cos(x+60)=\cos(x)\cos(60)-\sin(x)\sin(60)\] now x is a variable, but 60 is a number, so evaluating we get \[\frac{1}{2}\cos(x)-\frac{\sqrt{3}}{2}\sin(x)=2\sin(x)\]

OpenStudy (anonymous):

so far so good?

OpenStudy (anonymous):

let me see.

OpenStudy (anonymous):

Ok, I'm fine up until the very last step. How did you get that?

myininaya (myininaya):

are you talking about how did he know \[\cos(60)=\frac{1}{2} ; \sin(60)=\frac{\sqrt{3}}{2}?\]

OpenStudy (anonymous):

\[\cos(60)=\frac{1}{2}\] and \[\sin(60)=\frac{\sqrt{3}}{2}\]

OpenStudy (anonymous):

i evaluated the funtions at 60 degrees

OpenStudy (anonymous):

Oh, Ok.

myininaya (myininaya):

We can actually proved this by drawing an equilateral triagle :)

OpenStudy (anonymous):

Ok, so next step?

OpenStudy (anonymous):

then i multiplied everything by 2 (both sides) to get \[\cos(x)-\sqrt{3}\sin(x)=4\sin(x)\]

OpenStudy (anonymous):

myininiaya is a triangle junkie. likes them almost as much as kit kats

OpenStudy (anonymous):

Triangle jukie? is this a new kind of addition?

OpenStudy (anonymous):

then put cosine by itself on the left and got \[\cos(x)=4\sin(x)+\sqrt{3}\sin(x)\] factored and got \[\cos(x)=(4+\sqrt{3})\sin(x)\]

OpenStudy (anonymous):

it is an addiction addition

OpenStudy (anonymous):

then divided by sine to get \[\frac{\cos(x)}{\sin(x)}=4+\sqrt{3}\]

OpenStudy (anonymous):

Ah Ok. Triangles... Wish I was a junkie in them! Ok, next! :)

myininaya (myininaya):

\[\cot(x)=4+\sqrt{3} \]

OpenStudy (anonymous):

then took the reciprocal and got \[\tan(x)=\frac{1}{4+\sqrt{3}}\]

OpenStudy (anonymous):

because my calculator does not have a cotangent button on it

myininaya (myininaya):

lol okay satellite

myininaya (myininaya):

so we are looking for approximations not exact solutions?

OpenStudy (anonymous):

and finally \[x=\tan^{-1}\left( \frac{1}{4+\sqrt{3}}\right)\]

OpenStudy (anonymous):

@myininaya yes i guess so, since i can't think of another way

OpenStudy (anonymous):

i thought maybe there was some trick to get an exact answer, but i can't see it

OpenStudy (anonymous):

I think that's the only way to get it. Thanks for the help! I understand it now! :)

myininaya (myininaya):

for more solutions: \[\tan(x+180 n)=4+\sqrt{3}\]

OpenStudy (anonymous):

I don't know if I could get the answer w/o the steps first being placed.... Must get the hang of it!

myininaya (myininaya):

where n is an integer

myininaya (myininaya):

\[x=\tan^{-1}(4+\sqrt{3})-180 n , n \in \mathbb{Z} \]

OpenStudy (anonymous):

what myininaya said. you are asked for all solutions in [0, 360)

myininaya (myininaya):

oh didn't see the interval

myininaya (myininaya):

but still you can use the formula i have to find all the solutions in that interval

OpenStudy (anonymous):

actually it is \[x=\tan^{-1}(\frac{1}{4+\sqrt{3}})-180 n , n \in \mathbb{Z}\]

OpenStudy (anonymous):

or something like that

myininaya (myininaya):

yes satellite is right

OpenStudy (anonymous):

Hmm. is that part of the answer?

OpenStudy (anonymous):

2 to 2200

OpenStudy (anonymous):

you get 9.896° so that is one answer

OpenStudy (anonymous):

then add 180 degrees for the other answer

OpenStudy (anonymous):

Ok. Thanks!